Example #5: What is the solubility of Ca(OH)2 in 0.0860 M Ba(OH)2?
Answer #1

Ba(OH)2-------------> Ba^2+ (aq) + 2OH^- (aq)

0.086M                                            2*0.086M

                                                        0.172M

                                                          

Ca(OH)2(s) ----------------> Ca^2+ (aq) + 2OH^- (aq)

                                         x                      2x+0.172

Ksp    =   [Ca^2+][OH^-]^2

4.68*10^-6   = x*(2x+0.172)^2

4.68*10^-6    = x*(0.172)^2                   [ 2x+0.172 = 0.172 , 2x<<<<<0.172]

x                =   0.000158

The solubility of Ca(OH)2 is 1.58*10^-4 M >>>>answer